A filler pipe can fill a tank in 10 hours, while an outlet pipe can empty the tank in 15 hours. How long will it take to fill the empty tank with both pipes operating?
Found 2 solutions by Simnepi, Alan3354:
Answer by Simnepi(216)

The tank will never be filled if both pipes operate at the same time.
this is because the outlet pipe can flow faster than the inlet pipe.

Answer by Alan3354(69443)

A filler pipe can fill a tank in 10 hours, while an outlet pipe can empty the tank in 15 hours. How long will it take to fill the empty tank with both pipes operating?
---------------
In one hour, the filler adds 1/10 of the tank.
The outlet releases 1/15 of the tank.
-----
Net gain per hour = 1/10 - 1/15 = 1/30 of the tank
--> 30 hours